OpenStudy (anonymous):

Can anyone help on this problem? (U)=100 (B')=48, (A n B')=26, (A n B)= 24 Find (A u B)?_______________________

OpenStudy (mathmate):

Assuming (A) = cardinality of A = no. of elements in A, etc.  (B')=48 => (B)=100-48=52 (AUB)=(A)+(B)-(AnB) [all are known except (AUB)